OpenStudy (anonymous):

H5P1 ZERO-OFFSET AMPLIFIER In many amplifiers, we use dual power supplies so we can obtain a 0V offset at the output, that is, the DC operating point at the output is 0V. An example amplifier circuit that can achieve this is shown below. In this circuit, VS+=1.0V, VS−=−1.0V, and the MOSFET parameters are K=1 mAV2 and VT=0.5V. what formulas do i need

OpenStudy (anonymous):

did you get correct the Max VIN?

OpenStudy (anonymous):

not yet. i'm still on it. can you help, please?

OpenStudy (anonymous):

Actually in finding valid Operating ranges \[V0 = [-1 + \sqrt{1+2*K*RL*Vs} ] \div KRL\] In the above formula Vs= applied voltage across RL and Drain to Source( i.e Vds).. \[Vs=ids∗RL+Vds(or)Vs=ids∗RL+V0\] But in given problem , by applying KVL to output loop we get \[Vds+idsRL=Vs+(−Vs−)\] \[Vds+ids∗RL=1−(−1)=2\] So in finding V0 substitute\[Vs=2 \] (NOT 1) And we know Valid maximum input range is \[Vgs = V0 + Vthershold\] Applying KVL to input loop \[Vin=Vgs+Vs−\] Substitue the value ...You will get it !!!!!! Cheers!!!!